# SpikeSMOKE: Spiking Neural Networks for Monocular 3D Object Detection with Cross-Scale Gated Coding

### v2

@@ Abstract
Low energy consumption for 3D object detection is an important research area because of the increasing energy consumption with their wide application in fields such as autonomous driving. The spiking neural networks (SNNs) with low-power consumption characteristics can provide a novel solution for this research. Therefore, we apply SNNs to monocular 3D object detection and propose the SpikeSMOKE architecture in this paper, which is a new attempt for low-power monocular 3D object detection. As we all know, discrete signals of SNNs will generate information loss and limit their feature expression ability compared with the artificial neural networks (ANNs). In order to address this issue, inspired by the filtering mechanism of biological neuronal synapses, we propose a cross-scale gated coding mechanism(CSGC), which can enhance feature representation by combining cross-scale fusion of attentional methods and gated filtering mechanisms. In addition, to reduce the computation and increase the speed of training, we present a novel light-weight residual block that can maintain spiking computing paradigm and the highest possible detection performance. Compared to the baseline SpikeSMOKE under the 3D Object Detection, the proposed SpikeSMOKE with CSGC can achieve 11.78 (+2.82, Easy), 10.69 (+3.2, Moderate), and 10.48 (+3.17, Hard) on the KITTI autonomous driving dataset by $AP|_{R_{11}}$ at 0.7 IoU threshold, respectively. It is important to note that the results of SpikeSMOKE can significantly reduce energy consumption compared to the results on SMOKE. For example, the energy consumption can be reduced by 72.2% on the hard category, while the detection performance is reduced by only 4%. SpikeSMOKE-L (lightweight) can further reduce the amount of parameters by 3 times and computation by 10 times compared to SMOKE.
@@ I Introduction
Although ANNs have shown excellent performance in various fields including computer vision (CV)[1], speech recognition (SR)[2], natural language processing (NLP)[3], etc, they are facing with skyrocketing problems of energy consumption, with the increasingly complex models and progressively large volumes of data. To solve this problem, some researchers are turning to brain-inspired SNNs, because they have excellent characteristics of event-driven computation[4], biologically interpretable[5], and asynchronous temporal processing.
@@
At present, there has been a great deal of excellent research on SNNs applied to computer vision (CV), natural language processing (NLP) and other fields[6][7][8][9][10][11][5][12]. For instance, advancements from Spiking ResNet[13] to EMS-ResNet[11] and then to MS-ResNet[12] introduced increasingly efficient event-driven network architectures. Spikformer[14], which is the initial utilization of Transformer in SNNs, followed by ongoing enhancements, led to the introduction of Spike-Driven Transformer v1/v2[15][16]. This innovation offers robust sequence modeling capabilities and heightened parallel processing efficiency for intricate tasks. For object detection, current research focuses on 2D object detection using 2D image data or 3D object detection using 3D data. For example, in Spiking-PointNet [17], point clouds—which represent 3D spatial information—were utilized for the purpose of detecting 3D objects, marking the first successful application of point clouds within SNNs. In [18], a spike-driven object detection framework SpikeYOLO was proposed for 2D image data. However, due to the lack of depth information, monocular 3D object detection using 2D image data for 3D object detection has not yet been explored through SNNs, which is unable to utilize the low-power characteristics of SNNs.
@@
Monocular 3D object detection combines 2D image information and depth estimation to achieve detection and localization of 3D objects, which has been widely used in the field of autonomous driving[19][20][21][22][23]. Currently, there are many excellent models in the field of ANNs. For example, MonoWAD [24] introduced an innovative weather-adaptive diffusion model for monocular 3D object detection, LabelDistill[25] presented label-guided cross-modal knowledge distillation for camera-based 3D object detection, Mono3DVG[26] offered a fresh perspective by detecting 3D objects through detailed descriptions of visual appearance, and MonoCD[27], a network architecture based on auxiliary information. However, when applying monocular 3D object detection to resource-constrained edge devices like autonomous driving[28], its practical deployment will be restricted. So it is very important to implement a network model with low-power characteristics for monocular 3D object detection. Naturally, brain-like SNNs model with low-power characteristics, can offer us a novel insight to solve this problem.
@@
Based on the above analysis, we apply SNNs to monocular 3D object detection to construct a novel architecture SpikeSMOKE. In order to shrink the gap between the discrete signals of SNNs and the continuous signals of ANNs for feature expressiveness, we propose a cross-scale gated coding mechanism (CSGC) inspired by the filtering mechanism of biological neuronal synapses, which can reduce information loss and improve the detection performance of the model by fusing multi-scale attention, as shown in Figure 1. To further reduce the model’s power consumption while maintaining the spike computation paradigm, we propose innovative light-weight residual blocks. Numerous experiments have demonstrated on the autonomous driving KITTI dataset and the CIFAR10/100 datasets. The contributions of this paper can be summarized as follows.
@@
(1) To enhance the energy efficiency of monocular 3D object detection, we draw inspiration from brain-like SNNs, known for their low-power properties, and leverage the simplicity of the SMOKE network to construct a novel architecture called SpikeSMOKE.
@@
(2) In this SpikeSMOKE, we introduce the parallel cross-scale gated coding mechanism CSGC based on attention to improve the information representation capability of the model. At the same time, we also propose a light-weight residual block on the SpikeSMOKE model.
(3) The SpikeSMOKE can significantly decrease energy consumption compared to the results with SMOKE for 3D Object Detection. For instance, energy usage can be reduced by 72.2% in the hard category, with merely a 4% dip in detection performance.
(4)The SpikeSMOKE with CSGC can achieve better results on the KITTI dataset, in comparison to the baseline SpikeSMOKE. Moreover, in order to validate the generalizability of the proposed CSGC encoding strategy, we have also done experiments on the CIFAR-10/100 datasets to validate this.
@@ II-A Spike Neuron
Spike neurons simulate the behavior of biological neurons transmitting signal sequences through synapses, transmitting temporal signals and spatial information through the membrane voltage U. The commonly used neurons in spiking neural networks consist of integrate-and-fire (IF)[29] neurons and leaky integrate-and-fire (LIF) [30]neurons. We used LIF neurons and described their neuronal dynamics formulation as follows:
@@
$U_{i}^{l}[t]$ denotes the membrane voltage of neuron i in layer l, $W_{ij}$ indicates the connection weights of layer l and layer l-1, $S_{j}^{l}[t]$ represents the spike of the jth neuron in layer l-1, and $Hea$ is the Heaviside function. At discrete time step T, when the membrane voltage of the neuron is greater than the spike neuron threshold, i. e. , $U_{i}^{l}[t]$$>$$U_{th}$, the neuron emit a spike[31]. $\tau$ denotes the decay coefficient. After issuing spike 1, the membrane voltage is usually reset to 0.
@@ II-B Monocular 3D Object Detection
Common approaches for monocular 3D object detection include Two-stage methods (MonoLSS [32]), Single-stage anchor-based methods (GrooMeD-NMS [33]) and Single-stage anchor-free methods (SMOKE[34], MonoCD [27], MonoDGP [35]). Compared with the first method, which is prone to introducing 2D noise[34], and the second method, which usually requires complex data preprocessing and non-maximum suppression[36][37], the third method has the advantages of a simple model and high computational efficiency. In this paper, we use the SMOKE [34] architecture of Single-stage anchor-free methods as a base architecture, which both ignores the redundancy associated with 2D detection frameworks and does not require additional data. The SMOKE architecture continues the keypoint detection of centernet [38], which discards the traditional 2D detection module and retains only the 3D detection part, and improves the parameter convergence and detection accuracy by multi-step disentanglement. The 3D bounding box is obtained by predicting the object’s 3D projection center on the image plane along with attribute variables. The attribute regression of the 3D bounding box is decoupled into an 8-tuple $(\alpha_{x},\alpha_{y},\alpha_{z},\alpha_{l},\alpha_{w},\alpha_{h},sin\beta,cos\beta)$. These parameters are optimized by a loss function, which is transformed to obtain the true parameters $(x,y,z,w,h,l,\theta)$ used to construct the 3D bounding box.
@@ II-C Coding Mechanism
In a spiking neural network it is necessary to first encode the data into the form of spikes, there are several common ways of coding:rating coding [39] encodes information based on the average rate of neuron firing, phase coding [40] is expressed by the temporal position of the spike relative to a certain reference event, temporal coding [41], [42], [43] represents information based on the time interval and order of the spikes, and frequency coding involves the frequency at which neurons emit spikes over a specific time period. In addition to this, there are various complex coding methods ([44][45][46][47] ), and choosing an appropriate coding method is crucial for the representation performance of the model, in order to reduce the gap between the discrete and continuous signals in the spiking neural network, we propose a novel gated coding method in conjunction with the cross-scale fusion unit.
@@ II-D Light-Weight Model
Light-weight model refers to reducing the number of parameters and computation and compressing the model training time by removing or transforming part of the model architecture and connections. There are several common methods for lightweighting models in neural networks, a) Pruning: Chen et al [48] proposed an SNN pruning regeneration mechanism based on neuron criticality to remove unnecessary connections or neurons in the network. SSNN [49] designed sparse RSNN by pruning randomly initialized model. b) Model quantization: Spiking-Diffusion [50] proposed a diffusion model based on vector quantization discretization. c) Model distillation: TSSD [51] proposed a spatio-temporal distillation method. d) Depth-wise separable convolutions: the first proposed in MobileNets [52], and later extended and applied widely in the field of ANN ( [53], [54], [55], [56]). There has been a great deal of experimental or theoretical evidence for the effectiveness of depth-wise separable convolutions, and our proposed lightweight residual unit originates from this idea.
@@ III-A SpikeSMOKE
SpikeSMOKE architecture, as shown in Figure 2, takes the single-stage object detection network SMOKE as a infrastructure and retains its macro-architecture, because this architecture can ignore the redundancy of 2D detection framework and does not require additional data.
Backbone: Since DLA34 can take advantage of fusing feature maps of different scales through the depth convergence feature, we use the spike firing rate of the spike neurons to mimic its ReLU activation function, converting it to Spike-DLA34 as our backbone.
Neck: In order to ensure that the model is fully spike-driven, we add a spike neuron before each vanilla convolution in the DLAup structure, which can use deformable convolution to improve feature representation.
Head: This component consists two branches. One is keypoint classification used by heatmap, the other is 3D bounding box regression. This two branches can process the feature maps from the Neck network to obtain the 3D object detection results.
@@ III-B Cross-Scale Gated Coding (CSGC)
For ANN2SNN, the information loss is inevitable because the continuous signals are transformed to the discrete signals. In order to improve the feature expression capability, we propose a cross-scale fusion of attention coding unit named as CSGC. This method can utilize different convolution kernel to fuse cross-scale contextual information.
For CSGC, we design a parallel architecture including channel attention and spatial attention and a gate control unit. The input $x\in{R^{B\times C\times H\times W}}$ is firstly repeated at each time step to obtain information in the time dimension. Then, in the channel attention part, we use a Linear-ReLU-Linear structure to learn and update the weights in the channel dimension by the Equation (2):
@@
In the spatial attention part, we conduct feature extraction on the same feature map by different convolution kernel (Equation (3)). Small convolution kernels are employed to capture local subtle features for small objects, enabling the detection of small-sized objects. Conversely, larger convolution kernels are utilized to cover larger areas for detecting large objects and effectively capturing global information. We employ three different sizes of convolution kernels and assign dynamic weights to the feature maps after different convolutional processing by the learnable parameters $\alpha$, $\beta$ and $\gamma$, respectively. At the same time, we invoke the residual idea of [57] to connect the original feature map with the output part.
@@
$O(x)=Sig\left(SA(x)\odot CA(x)\right),$ (4)
where $\odot$ denotes the Hadamard product. Then, it performs the Hadamard product operation with the output of the spiking neuron, imitating the filtering function of the synapses, as defined by Equation (5):
$\widehat{O}(x)=O(x)\odot SNN(x).$ (5)

### v3

@@ Abstract
With the wide application of 3D object detection in some fields such as autonomous driving, its energy consumption is constantly increasing, making the research on low-power consumption alternatives a key research area. The spiking neural networks (SNNs), possessing low-power consumption characteristics, offer a novel solution for this research. Consequently, we apply SNNs to monocular 3D object detection and propose the SpikeSMOKE architecture, which represents a new attempt at low-power monocular 3D object detection. It’s well known that the discrete signals of SNNs can lead to information loss compared to artificial neural networks (ANNs), which limits their feature representation capabilities. To solve this problem, inspired by the synaptic filtering mechanism of biological neurons, we propose a new Cross-Scale Gating Coding Mechanism (CSGC), which can enhance feature representation by combining cross-scale fusion of attentional methods and gated filtering mechanisms. In addition, to reduce the computation and accelerate training, we present a novel light-weight residual block that can maintain spiking computing paradigm and the highest possible detection performance. Our method is effective on the KITTI, NuScenes-mini and CIFAR10/100 datasets. Compared to the baseline SpikeSMOKE under the 3D Object Detection, the proposed SpikeSMOKE with CSGC can achieve 11.78 (+2.82, Easy), 10.69 (+3.2, Moderate), and 10.48 (+3.17, Hard) on the KITTI autonomous driving dataset by $AP|_{R_{11}}$ at 0.7 IoU threshold, respectively. It is worth noting that the results of SpikeSMOKE can significantly reduce energy consumption compared with the results of SMOKE. And SpikeSMOKE-L (lightweight) can further reduce the amount of parameters by 3 times and computation by 10 times compared to SMOKE.
@@ I Introduction
Although ANNs have shown excellent performance in various fields including computer vision (CV)[22], speech recognition (SR)[11], natural language processing (NLP)[19], etc, and have made great contributions to the development of artificial intelligence, they are facing with skyrocketing problems of energy consumption, with the increasingly complex models and progressively large volumes of data[56]. Especially in 3D object detection, the maximum computational complexity observed in existing algorithms has already reached 50 GMAC[28]. This not only increases hardware costs but also limits their applications in mobile devices and edge computing scenarios. To solve this problem, some researchers are turning to brain-inspired SNNs, because they have excellent characteristics of event-driven computation[61], biologically interpretable[15], and asynchronous temporal processing. These features enable SNNs to significantly reduce energy consumption and improve computational efficiency when dealing with sparse data and temporal dynamic signals.
@@
At present, there has been a great deal of excellent research on SNNs applied to computer vision (CV), natural language processing (NLP) and other fields[53][50][35][8][58][16]. In terms of the evolution of network architecture, the early Spiking ResNet[57] laid the foundation for event-driven network architectures. Subsequently, EMS-ResNet[9] optimized this foundation, and MS-ResNet[16] further enhanced the network’s ability to process complex data, bringing new ideas to the field of computer vision. In the realm of natural language processing, Spikformer[60], which is the initial utilization of Transformer in SNNs, opened a new path for sequence modeling. The subsequent Spike-Driven Transformer v1/v2[49][48] continuously optimized, which not only offers robust sequence modeling capabilities, but also improves the efficiency of parallel processing of intricate language tasks. For object detection, current research focuses on 2D object detection using 2D image data or 3D object detection using 3D data. For example, in Spiking-PointNet [38], point clouds—which represent 3D spatial information—were utilized for the purpose of detecting 3D objects, marking the first successful application of point clouds within SNNs. In aspect of 2D image data, SpikeYOLO[31] has put forward a spike-driven object detection framework, which is capable of effectively extracting target features from 2D images and carrying out target localization and classification. However, SNNs still face some challenges in the field of object detection, especially in monocular 3D object detection. Since monocular 3D object detection requires processing additional depth information, traditional 2D object detection methods cannot be directly applied to 3D object detection. Up to now, there has been no mature exploration of monocular 3D object detection via SNNs, and the low-power characteristic of SNNs has not been utilized in this task.
@@
Monocular 3D object detection combines 2D image information and depth estimation to achieve detection and localization of 3D objects, which has been widely used in the field of autonomous driving[27][47][10][18]. Currently, there are many excellent models in the field of ANNs. For example, MonoWAD [33] introduced an innovative weather-adaptive diffusion mechanism, which can effectively deal with changes in image features under different weather conditions, thereby enhancing the robustness and accuracy of monocular 3D object detection. LabelDistill[21] presented a label-guided cross-modal knowledge distillation method, which uses data from other modalities to assist in 3D object detection from 2D images. Mono3DVG[52] offered a fresh perspective by detecting 3D objects through detailed descriptions of visual appearance, and is able to accurately identify and locate objects with complex appearances. MonoCD[46], a network architecture based on auxiliary information, further optimized the performance of monocular 3D object detection. However, when applying monocular 3D object detection to resource-constrained edge devices like autonomous driving[25], its practical deployment will be restricted. These edge devices typically have limited computing resources and energy supplies, while existing ANN models often require high computational power and energy consumption to implement complex detection algorithms. Therefore, it is very important to implement a network model with low-power characteristics for monocular 3D object detection. Against this backdrop, brain-inspired SNNs models have attracted attention for their unique low-power characteristics, which give them potential advantages in handling monocular 3D object detection tasks and offer new ideas for solving this problem.
@@
Based on the above analysis, we apply brain-inspired Spiking Neural Networks (SNNs) with low-power characteristics to monocular 3D object detection to construct a novel architecture SpikeSMOKE. In order to shrink the gap between the discrete signals of SNNs and the continuous signals of ANNs for feature expressiveness, we propose a cross-scale gated coding mechanism (CSGC) inspired by the filtering mechanism of biological neuronal synapses, which can reduce information loss and improve the detection performance of the model by fusing multi-scale attention. To further reduce the model’s power consumption while maintaining the spike computation paradigm, we design innovative light-weight residual blocks that can significantly reduce the computational load and energy consumption while maintaining the network’s depth and complexity. To verify the effectiveness of the proposed method in this paper, we conducted a large number of experiments, which cover the widely used KITTI and NuScenes-mini datasets in the field of autonomous driving, as well as the CIFAR10 and CIFAR100 datasets in the field of computer vision. The contributions of this paper can be summarized as follows.
@@
(1) To enhance the energy efficiency of monocular 3D object detection, we propose a spiking architecture SpikeSMOKE based on SMOKE leveraging the low power consumption feature of brain-like SNNs.
@@
(2) To improve the information representation capability of SpikeSMOKE, we design a novel parallel cross-scale gated coding mechanism CSGC based on attention. Additionally, we propose a light-weight residual block integrated into the SpikeSMOKE model.
(3) Compared to the baseline SpikeSMOKE, our enhanced model with CSGC achieves superior performance on the KITTI dataset. We further validate the generalizability of the proposed CSGC encoding strategy through experiments on the NuScenes-mini and CIFAR-10/100 datasets. Meanwhile, compared with SMOKE, SpikeSMOKE significantly reduces energy consumption in 3D object detection.
@@ II-A Spike Neuron
Spike neurons simulate the behavior of biological neurons transmitting signal sequences through synapses[43], transmitting temporal signals and spatial information through the membrane voltage U. This mechanism enables spike neurons to efficiently process dynamic information while maintaining low energy consumption, which is particularly significant when dealing with sparse data and time-series tasks. The commonly used neurons in spiking neural networks consist of integrate-and-fire (IF)[32] neurons and leaky integrate-and-fire (LIF) [40][17]neurons. These neuron models simulate the signal processing of biological neurons through different mechanisms[24]. We used LIF neurons and described their neuronal dynamics formulation as follows:
@@
[…]
@@ II-B Monocular 3D Object Detection
Monocular 3D object detection relies solely on the 2D image information obtained from a monocular camera to detect and locate 3D objects in a scene. However, the lack of depth information in monocular images makes it highly challenging to recover 3D information from 2D images. To address this challenge, researchers have proposed a variety of methods for monocular 3D object detection, which can generally be divided into three categories: two-stage methods, single-stage anchor-based methods, and single-stage anchor-free methods. Two-stage methods, such as MonoLSS[26], typically first generate a series of candidate regions and then perform fine classification and regression on these regions to determine the final 3D object bounding boxes. This method is prone to introducing 2D noise[30], which to some extent affects the accuracy and reliability of detection. Single-stage anchor-based methods, such as GrooMeD-NMS[23], define a series of anchor points on the image in advance and then perform classification and regression on these anchor points to directly generate 3D bounding boxes. These methods usually require complex data preprocessing and non-maximum suppression[6][36]. Compared with the first two categories of methods, single-stage anchor-free methods (SMOKE[30], MonoCD [46], MonoDGP [34]) have the significant advantages of a simple model and high computational efficiency, making them more suitable for real-time operation on resource-constrained edge devices.
In this paper, we use the SMOKE [30] architecture of Single-stage anchor-free methods as a base architecture, which both ignores the redundancy associated with 2D detection frameworks and does not require additional data. The SMOKE architecture continues the keypoint detection of centernet [7], which discards the traditional 2D detection module and retains only the 3D detection part, and improves the parameter convergence and detection accuracy by multi-step disentanglement. Specifically, it decomposes the attribute regression of the 3D bounding box into multiple steps, optimizing each parameter progressively. The 3D bounding box is obtained by predicting the object’s 3D projection center on the image plane along with attribute variables. The attribute regression of the 3D bounding box is decoupled into an 8-tuple $(\alpha_{x},\alpha_{y},\alpha_{z},\alpha_{l},\alpha_{w},\alpha_{h},sin\beta,cos\beta)$. These parameters are optimized by a loss function, and then transformed into the actual parameters for constructing the 3D bounding box, including the object’s center position $(x,y,z)$, dimensions $(w,h,l)$, and orientation angle $(\theta)$.
@@ II-C Coding Mechanism
In a spiking neural network, data encoding is a crucial step for efficient information processing. Initially, input data must be transformed into spike sequences for further processing by neurons, a process that involves selecting from several encoding methods. Rating coding [42] encodes information based on the average rate of neuron firing, the higher the firing rate of neurons, the greater the signal strength or information content they represent. Its advantage lies in its simple implementation and similarity to the behavior of biological neurons. However, it requires a longer time window to accurately estimate the firing rate, which may lead to information delay. Phase coding [20] expresses information by the temporal position of spikes relative to a reference event. For instance, in a periodic stimulus, the phase of a spike can precisely indicate specific signal features. Temporal coding [59] represents information based on the time intervals and order of spikes, effectively utilizing the temporal characteristics of spikes to handle dynamic and sequential data. Frequency coding involves the frequency at which neurons emit spikes over a specific time period. It represents information through the repetition rate of spikes and emphasizes short-term frequency changes rather than long-term average firing rates. In addition to this, there are various complex coding methods ([37][45] ), and choosing an appropriate coding method is crucial for the representation performance of the model. In order to reduce the gap between the discrete and continuous signals in the spiking neural network, we propose a novel gated coding method in conjunction with the cross-scale fusion unit. This encoding method introduces a gating mechanism that more effectively controls the flow and processing of information.
@@ II-D Light-Weight Model
Light-weight model refers to reducing the number of parameters and computation and compressing the model training time by removing or transforming part of the model architecture and connections[44][51]. There are several common methods for lightweighting models in neural networks, a) Pruning: Chen et al [4] proposed an SNN pruning regeneration mechanism based on neuron criticality to remove unnecessary connections or neurons in the network. This mechanism can identify and remove connections or neurons that contribute less to the model’s performance, thereby simplifying the network structure and reducing unnecessary computations. SSNN [2] designed sparse RSNN by pruning randomly initialized model. b) Model quantization: Model quantization reduces the model’s storage requirements and computational load significantly by converting the model’s weights and activations from floating-point numbers to low-bit representations, such as binary or ternary quantization. Spiking-Diffusion [29] proposed a diffusion model based on vector quantization discretization. c) Model distillation: TSSD [62] proposed a spatio-temporal distillation method, which transfers the knowledge from a complex large model (the teacher model) to a simpler small model (the student model). This enables the student model to inherit the performance of the teacher model while maintaining low computational complexity. d) Depth-wise separable convolutions: Depth-wise separable convolutions decompose standard convolutions into two steps: depth-wise convolutions and point-wise convolutions, significantly reducing the computational load of convolution operations. The first proposed in MobileNets [14], and later extended and applied widely in the field of ANN ( [5], [3]). There has been a great deal of experimental or theoretical evidence for the effectiveness of depth-wise separable convolutions, and our proposed lightweight residual unit originates from this idea, aiming to further optimize the model’s computational efficiency and performance.
@@ III-A SpikeSMOKE
The SpikeSMOKE architecture is shown in Figure 1. The design of this architecture is based on the single-stage object detection network SMOKE, which has the advantages of ignoring the redundancy of 2D detection frameworks and not requiring additional data. We choose this architecture because it can simplify the complexity of the model while maintaining efficient detection performance.
Backbone: The architecture DLA34 can take advantage of fusing feature maps of different scales through the depth convergence feature. To fully transform the model into a SNNs-based architecture, we introduced the spiking rate of spiking neurons to replace the ReLU activation function in DLA34, which is very crucial. Meanwhile, to enhance the processing capability of spike signals, we have proposed the CSGC module in backbone, including channel attention mechanism and spatial attention mechanism. A detailed introduction to CSGC is in Section Methods B.
Neck: The design of the neck aims to further optimize the transmission and processing of features, ensuring that the model is fully driven by spikes. In the DLAup structure, we ensure that a spiking neuron is used before each convolution operation to convert the continuous signals into discrete spike signals, thereby maintaining the spike-driven characteristics. Meanwhile, deformable convolutions were employed to further enhance feature extraction.
Head: This component consists two branches. One is keypoint classification used by heatmap, which is mainly responsible for analyzing the input feature maps to identify the locations of object keypoints and generate heatmaps to represent the distribution of these keypoints. The other is 3D bounding box regression, which is in charge of predicting the 3D bounding box parameters of each detected object based on the information in the feature maps, including the center coordinates, size, and rotation angles of the bounding boxes. This two branches can process the feature maps from the Neck network to obtain the 3D object detection results.
@@ III-B Cross-Scale Gated Coding (CSGC)
When converting ANN to SNN, the fundamental contradiction of sparsity and discreteness of SNN is faced, which weakens the modeling ability of continuous spatial structures. Especially in monocular 3D detection tasks, the model needs to recover nonlinear 3D geometric information from 2D images, which poses high requirements for the multi-scale context awareness ability and key structure retention ability of features. However, standard SNN encoders (such as direct ANN2SNN conversion) lose a large amount of information during the spiking process. Therefore, in order to enhance the feature expression ability and reduce information loss, we propose a cross-scale fusion of attention coding unit named as CSGC, as shown in Figure 2.
For CSGC, we design a parallel architecture including channel attention and spatial attention and a gate control unit. The Channel Attention addresses the issue of which feature channels are more important for 3D geometric reasoning. The Spatial Attention is used to focus on which areas in an image are more critical to the current target scale. The original input tensor $x\in{R^{C\times H\times W}}$ is firstly repeated at each time step to obtain information in the time dimension. Then, the original tensor is fed into both the channel attention and spatial attention modules. In the channel attention part, we use a Linear-ReLU-Linear structure to learn and update the weights in the channel dimension by the Equation (2):
@@
The number of neurons in the middle layer of the network is set to the number of input channels /ratio, and the ratio is set to 4. This structure can effectively highlight important channel features, thereby enhancing the model’s ability to perceive key information. In the spatial attention part, we conduct feature extraction on the same feature map by different convolution kernel (Equation (3)). Small convolution kernels are employed to capture local subtle features for small objects, enabling the detection of small-sized objects. Conversely, larger convolution kernels are utilized to cover larger areas for detecting large objects and effectively capturing global information. To better integrate information from different scales, we employ three different sizes of convolution kernels and assign dynamic weights to the feature maps after different convolutional processing by the learnable parameters $\alpha$, $\beta$ and $\gamma$, respectively, with the initial weights all set to 0.5. At the same time, we invoke the residual idea of [12] to connect the original feature map with the output part. This not only enhances the features but also prevents information loss, ensuring that the model can fully utilize all useful information from the original input.
@@
$O(x)=\sigma\left(SA(x)\odot CA(x)\right),$ (4)
where $\sigma$ denotes the Sigmoid function, which maps values to the range (0,1), and $\odot$ denotes the Hadamard product. Then, it performs the Hadamard product operation with the output of the spiking neuron, imitating the filtering function of the synapses, as defined by Equation (5):
$\widehat{O}(x)=O(x)\odot SNN(x),$ (5)
where $O(x)$ represents the attention scores obtained after the parallel channel attention and spatial attention modules, indicating the importance of the detected features. $SNN(x)$ denotes the output spikes from the original feature x after processing through the Leaky Integrate-and-Fire (LIF) neuron. Since in SNNS, features exist in the form of binary pulses, direct weighting would undermine their sparsity and event-driven characteristics. Therefore, CSGC takes the attention score $O(x)$ as the gate signal and performs a Hadamard product with the binary output (0 or 1) of $SNN(x)$, which is generated by an LIF neuron, allowing only the spikes in the high-attention region to pass through, thereby obtaining the final output $\widehat{O}(x)$. This essentially simulates the filtering mechanism in biological synapses, achieving feature screening while maintaining the sparsity of spikes.